# The Phase Space Distance Between Collider Events

### v1

@@ 1 Introduction
Novel representations of collider data play an increasingly important role in the modern machine learning (ML) era Larkoski:2017jix ; Guest:2018yhq ; Albertsson:2018maf ; Radovic:2018dip ; Carleo:2019ptp ; Bourilkov:2019yoi ; Schwartz:2021ftp ; Karagiorgi:2021ngt ; Shanahan:2022ifi . In particular, a physically-motivated notion of the distance between collider events is a powerful input to ML-based methods for event classification and anomaly detection. Recently, optimal transport distances between energy flow distributions Komiske:2019fks ; Komiske_2020 ; PhysRevD.105.076003 ; Cai:2020vzx or spectral functions Larkoski:2023qnv have given rise to a new geometry of collider events, with widespread applications Komiske:2019jim ; Cesarotti:2020hwb ; Cesarotti:2020ngq ; Stein:2020rou ; Kasieczka:2021xcg ; Howard:2021pos ; DiGuglielmo:2021ide ; Kansal:2021cqp ; Collins:2021pld ; Cheng:2020dal ; CrispimRomao:2020ejk ; ATLAS:2023mny ; Craig:2024rlv . However, there is another natural distance between collider events that has yet to be fully explored: the one furnished by the phase space manifold on which scattering amplitudes are defined.
@@ 2.1 Phase Space in the CM Frame: A Product Manifold
Here $Q=(Q,0,0,0)$ is the total energy of the event in the CM frame and $\delta^{+}(p_{k}^{2})=\delta(p_{k}^{2})\,\Theta(p_{0k})$ enforcing that all particles are on-shell with positive energy. The on-shell constraints can be satisfied by re-expressing the four-momenta in terms of the two spinors for each particle, where
@@
in the CM frame. Notice that $\vec{u}$ is a real vector while $\vec{v}$ is complex. Written in terms of these $\vec{u},\vec{v}$ vectors, Eq. (1) now becomes
$\begin{split}d\Pi_{N}=&(2\pi)^{4-3N}Q^{2N-4}\left[\prod_{k=1}^{N}\frac{d^{2}\lambda_{k}^{1}d^{2}\lambda_{k}^{2}}{\text{U}(1)}\right]\\ =&(2\pi)^{4-3N}Q^{2N-4}\,\frac{d^{N}u\,d^{N}v}{\text{U}(1)^{N}}\,\delta(1-|\vec{u}|^{2})\,\delta(1-|\vec{v}|^{2})\,\delta^{(2)}(\vec{u}^{\dagger}\vec{v})\,,\end{split}$ (5)
@@
One can easily check that dimensionality works since the sum of the simplex and sphere dimensions gives us $(N-1)+(2N-3)=3N-4$, which is the correct phase space dimension for a system of $N$ massless particles. Also, because we will always compare events with the same total energy $Q$, we will often rescale energies so that $Q=1$, which eliminates the overall dimensionality of the phase space metric.
@@ 2.2 Phase Space in the CM Frame: Metric and Explicit Coordination
In this paper, we are more interested in the distances between pairs of events assigned according to their positions on the phase space manifold, so these line elements can be promoted to honest Riemannian metrics that satisfy symmetry, identity of indiscernibles, and the triangle inequality. For two events that we label $A$ and $B$, their metric distance on the simplex is
@@
Distances on the hypersphere are determined by the angle between the two events, and therefore the metric on the hypersphere is
@@
where $\Re{\vec{v}_{A}^{\prime\dagger}\vec{v}_{B}^{\prime}}$ is the real part of the inner product of the vectors $\vec{v}^{\prime}_{A}$ and $\vec{v}^{\prime}_{B}$. Any convex linear combination of these two metrics is itself a metric, and could therefore be a metric on the phase space manifold. However, the phase space volume element is fixed, which constrains the determinant of the metric.
By demanding that the metric on the phase space reproduces the dependence on multiplicity $N$ in the volume element of Eq. (10), we can express the phase space metric between two events $A$ and $B$ as dependent on a single parameter $c>0$, where
@@ 2.3 Phase Space in the Collinear Limit
Note that unlike their definition for phase space in the CM frame, these vectors are dimensionless, which is preferred because to good approximation jets are scale invariant. Phase space volume in these coordinates then becomes
@@
in order to ensure that probabilities are dimensionless. Further, assuming scale invariance, probabilities are unchanged if the ratio
@@ 4.1.1 Results
Fig. 4 shows the classification performance of the total phase space distances coupled with SVM. The highest AUC obtained among all definitions is $\sim 0.86$, corresponding to the simplex definition of $\rho=p_{0}-p_{3}$; a relative weighting of sphere:simplex = 1:1; the $(p_{0}+p_{3})$-ordering; and an event representation of $N=4$ jets per event. This is quite remarkable, given that all we have utilized is the “locations” on the phase space manifold of different types of events. No explicit information about the underlying collision process is provided to the model.
Here we make three overall observations. First, as noted above, both the definitions of the simplex coordinates and the relative weighting between the simplex and the sphere make little difference on the final AUC score, suggesting that any valid phase space distance contains comparable information recoverable by SVM. Of course, this is only true within reasonable variation of the sphere-to-simplex weighting, as extreme weights asymptote to the sphere-only and simplex-only distances.
@@ 4.2.1 Results
Fig. 9 shows the AUC scores of the total phase space distances. The performance is excellent across a wide range of $N$ values, with higher $N$ giving an AUC score close to 1. This indicates that the current signals and backgrounds are extremely well separable on the phase space manifold, and that the higher the dimension of the phase space, the better the separability. Such observations correspond well with our physics intuition, since the new physics events have a much higher multiplicity than the QCD events.
We again note that both definitions of the $\vec{u}$ coordinates give comparable performance and the sphere-to-simplex relative weighting has little impact on classification, corroborating our hypothesis that any valid phase space distance definition contains equivalent information for a reasonable classifier to utilize. Furthermore, there is very little variation in the final performance for the three particle orderings, suggesting that the local minima found by the approximation schemes are likely very close to the global minimum.
@@ 5 Conclusions and Future Directions
There are a number of promising directions for future study. In converting explicit global coordinates on the phase space manifold to a practical metric, we have made certain choices that leave room for exploration. For example, as mentioned above, we have fixed the dimensionality of the phase space manifold for a given ensemble of events by exclusively clustering into $N$ objects, mapping events onto an $N$-object representation with crucial influence over the effectiveness of ML-based classification. Alternative approaches to treating events of different intrinsic dimensionality may prove fruitful.
Furthermore, the phase space distances computed between these $N$-object representations assume a massless phase space, which is not guaranteed to be a good approximation in all regimes. This could be circumvented by generalizing the phase space metric to massive particles. Finally, we forego an exact (yet computationally prohibitive) minimization of the distance over particle permutations in favor of an approximate procedure that is only guaranteed to find local minima. Alternative approaches to the minimization problem (potentially using machine learning) may prove effective.
Further opportunities also abound within our prescription for computing phase space distances. Analytic and numerical study of the phase space structure of different scattering amplitudes is likely to be productive. And while we have demonstrated the value of phase space distances in classification tasks, they are equally likely to be useful in other collider physics applications such as unsupervised anomaly detection and simulation-based inference.
@@ Acknowledgements.
We would like to thank Liam Brennan, Brian Henning, Joe Incandela, Tamas Vami, and Danyi Zhang for useful discussions. The work of TC, NC, and GK was supported in part by the U.S. Department of Energy under the grant DE-SC0011702 and performed in part at the Kavli Institute for Theoretical Physics, supported by the National Science Foundation under Grant No. NSF PHY-1748958. TC is additionally supported by the U.S. Department of Energy Awards No. DE-FOA-0002705, KA/OR55/22 (AIHEP) and No. DE-AC02-76SF00515.

### v2

# SLAC-PUB-17780

@@ 1 Introduction
Novel representations of collider data play an increasingly important role in the modern machine learning (ML) era Larkoski:2017jix ; Guest:2018yhq ; Albertsson:2018maf ; Radovic:2018dip ; Carleo:2019ptp ; Bourilkov:2019yoi ; Schwartz:2021ftp ; Karagiorgi:2021ngt ; Shanahan:2022ifi . In particular, a physically-motivated notion of the distance between collider events is a powerful input to ML-based methods for event classification and anomaly detection. Recently, optimal transport distances between energy flow distributions Komiske:2019fks ; Komiske_2020 ; PhysRevD.105.076003 ; Cai:2020vzx ; cai2021linearizedhellingerkantorovichdistance or spectral functions Larkoski:2023qnv have given rise to a new geometry of collider events, with widespread applications Komiske:2019jim ; Cesarotti:2020hwb ; Cesarotti:2020ngq ; Stein:2020rou ; Kasieczka:2021xcg ; Howard:2021pos ; DiGuglielmo:2021ide ; Kansal:2021cqp ; Collins:2021pld ; Cheng:2020dal ; CrispimRomao:2020ejk ; ATLAS:2023mny ; Craig:2024rlv ; Gouskos_2023 . However, there is another natural distance between collider events that has yet to be fully explored: the one furnished by the phase space manifold on which scattering amplitudes are defined.
@@ 2.1 Phase Space in the CM Frame: A Product Manifold
Here $Q=(Q,0,0,0)$ is the total energy of the event in the CM frame and $\delta^{+}(p_{k}^{2})=\delta(p_{k}^{2})\,\Theta(p_{0k})$ enforcing that all particles are on-shell with positive energy. Note that the phase space in Eq. (1) has a mass dimension of $42N-$ 4, whereas its number of degrees of freedom is $43N-$ 4. Unless otherwise specified, in our paper the term “dimension” refers to the number of degrees of freedom of a system, as this determines the dimension of the phase space manifold.
The on-shell constraints can be satisfied by re-expressing the four-momenta in terms of the two spinors for each particle, where
@@
in the CM frame. Notice that $\vec{u}$ is a real vector while $\vec{v}$ is complex, and both are dimensionless in the unit of mass. Written in terms of these $\vec{u},\vec{v}$ vectors, Eq. (1) now becomes
$\begin{split}d\Pi_{N}=&(2\pi)^{4-3N}Q^{2N-4}\,\frac{d^{N}u\,d^{N}v}{\text{U}(1)^{N}}\,\delta(1-|\vec{u}|^{2})\,\delta(1-|\vec{v}|^{2})\,\delta^{(2)}(\vec{u}^{\dagger}\vec{v})\,,\end{split}$ (5)
@@
One can easily check that dimensionality of the degrees of freedom works since the sum of the simplex and sphere dimensions gives us $(N-1)+(2N-3)=3N-4$, which is the correct phase space dimension for a system of $N$ massless particles. Also, because we will always compare events with the same total energy $Q$, we will often rescale energies so that $Q=1$, which eliminates the overall dimensionality of the phase space metric.
@@ 2.2 Phase Space in the CM Frame: Metric and Explicit Coordination
In this paper, we are more interested in the distances between pairs of events assigned according to their positions on the phase space manifold, so these line elements can be promoted to honest Riemannian metrics that satisfy symmetry, identity of indiscernibles, and the triangle inequality. For two events that we label $A$ and $B$, their metric distance on the simplex can be defined as
@@
Distances on the hypersphere are determined by the angle between the two events, and therefore a simple choice of the metric on the hypersphere is
@@
where $\Re{\vec{v}_{A}^{\prime\dagger}\vec{v}_{B}^{\prime}}$ is the real part of the inner product of the vectors $\vec{v}^{\prime}_{A}$ and $\vec{v}^{\prime}_{B}$.
Any convex linear combination of these two metrics is itself a metric, and could therefore be a metric on the phase space manifold. However, the phase space volume element is fixed, which constrains the determinant of the metric. By demanding that the metric on the phase space reproduces the dependence on multiplicity $N$ in the volume element of Eq. (10), we can express the phase space metric between two events $A$ and $B$ as dependent on a single parameter $c>0$, where
@@
In principle, one can further generalize the above definitions of the metrics on the simplex, the hypersphere, and their resulting product space, as long as the phase space volume element is preserved. This would give us freedom to modify the relative distance between events in different regions of the phase space. For example, one could define the distance such that more emphasis is put on the edge of the phase space in comparison to its bulk. Such generalized metrics may potentially lead to improved discriminative power than the ones defined above. However, we do not pursue this study here, in order not to unduly complicate our analysis. For us, our metrics on the simplex, sphere, and the phase space are simply defined by Eq. (16), Eq. (17), and Eq. (18), respectively.
@@ 2.3 Phase Space in the Collinear Limit
Again, these vectors are dimensionless in the mass unit, which is preferred because to good approximation jets are scale invariant. Phase space volume in these coordinates then becomes
@@
in order to ensure that probabilities are dimensionless in the mass unit. Further, assuming scale invariance, probabilities are unchanged if the ratio
@@ 4.1.1 Results
Fig. 4 shows the classification performance of the total phase space distances coupled with SVM. The highest AUC obtained among all definitions is $\sim 0.86$, corresponding to the simplex definition of $\rho=p_{0}-p_{3}$; a relative weighting of sphere:simplex = 1:1; the $(p_{0}+p_{3})$-ordering; and an event representation of $N=4$ jets per event. This is quite remarkable, given that all we have utilized is the “locations” on the phase space manifold of different types of events. No explicit information about the underlying collision process is provided to the model. In comparison, the best performing event shape observable studied in Ref. Cesarotti:2020hwb , i.e., the ring-like event isotropy, achieves an AUC score of only 0.774, about $10\%$ lower than our top performer. Of course, caution must be taken to interpret this comparison, given that our datasets are not the same and that Ref. Cesarotti:2020hwb employed only a simple cut on the single observable whereas here we use SVM.
We make three overall observations for the performance of the phase space metric. First, as noted above, both the definitions of the simplex coordinates and the relative weighting between the simplex and the sphere make little difference on the final AUC score, suggesting that any valid phase space distance contains comparable information recoverable by SVM. Of course, this is only true within reasonable variation of the sphere-to-simplex weighting, as extreme weights asymptote to the sphere-only and simplex-only distances.
@@ 4.2.1 Results
Fig. 9 shows the AUC scores of the total phase space distances. The performance is excellent across a wide range of $N$ values, with higher $N$ giving an AUC score close to 1 (i.e., 0.982). This indicates that the current signals and backgrounds are extremely well separable on the phase space manifold, and that the higher the dimension of the phase space, the better the separability. Such observations correspond well with our physics intuition, since the new physics events have a much higher multiplicity than the QCD events. Our near-perfect performance also echoes with the high AUC score of $\sim 0.930$ in Ref. Cesarotti:2020hwb , obtained again by the ring-like event isotropy. Both results indicate the relative ease of this particular classification problem.
For our phase space method, we again note that both definitions of the $\vec{u}$ coordinates give comparable performance and the sphere-to-simplex relative weighting has little impact on classification, corroborating our hypothesis that any valid phase space distance definition contains equivalent information for a reasonable classifier to utilize. Furthermore, there is very little variation in the final performance for the three particle orderings, suggesting that the local minima found by the approximation schemes are likely very close to the global minimum.
@@ 5 Conclusions and Future Directions
There are a number of promising directions for future study. In converting explicit global coordinates on the phase space manifold to a practical metric, we have made certain choices that leave room for exploration. For example, we forego an exact (yet computationally prohibitive) minimization of the distance over particle permutations in favor of an approximate procedure that is only guaranteed to find local minima. Alternative approaches to the minimization problem (potentially using machine learning) may prove effective.
We have also fixed the dimensionality of the phase space manifold for a given ensemble of events by exclusively clustering into $N$ objects, mapping events onto an $N$-object representation with crucial influence over the effectiveness of ML-based classification. Alternative approaches to treating events of different intrinsic dimensionality may prove fruitful, and one may also want to explore the impact of different clustering algorithms (such as inclusive clustering) on the classification performance of the phase space distance.
In this work, we have only studied the simplest choices of the metrics on the simplex, the hypersphere, and their product phase space. In principle, the metrics can be generalized to be functions of the coordinates, as long as they induce the expected topology and the phase space volume element is preserved. Such generalization would grant extra freedom to adjust the distance between events according to their place in the phase space, which may lead to even better classification performance. Furthermore, the computation of the phase space distances currently makes the assumption of massless objects, which is not guaranteed to be a good approximation in all regimes. This could be circumvented by generalizing the phase space metric to massive particles, and would make the current framework more versatile.
On the practical side, while we have demonstrated the value of phase space distances in classification tasks, they are equally likely to be useful in other collider physics applications such as unsupervised anomaly detection and simulation-based inference. Further analytic and numerical study of the phase space structure of different scattering amplitudes is most likely to be productive.
@@ Acknowledgements.
We would like to thank Liam Brennan, Brian Henning, Joe Incandela, Tamas Vami, and Danyi Zhang for useful discussions. We extend our special thanks to the referee, who suggested several generalizations of the framework for future study. The work of TC, NC, and GK was supported in part by the U.S. Department of Energy under the grant DE-SC0011702 and performed in part at the Kavli Institute for Theoretical Physics, supported by the National Science Foundation under Grant No. NSF PHY-1748958. TC is additionally supported by the U.S. Department of Energy Awards No. DE-FOA-0002705, KA/OR55/22 (AIHEP) and No. DE-AC02-76SF00515.